OpenStudy (shelby1290):

A bike ramp has an incline of 49 degrees. The height of the ramp is 0.8 metres. What is the distance the cyclist will travel?

OpenStudy (shelby1290):



OpenStudy (shelby1290):

This is what i've got so far... sin0 = opp/hyp sin49 / 1= 0.8/x x(sin49)=1(0.8) x(sin49)/sin49=0.8/sin49 x=1.06

OpenStudy (campbell_st):

your working and answer is correct

OpenStudy (shelby1290):

thank you!